A father who is 42 years old has a son and a daughter. The daughter is three times as old as the son. In 10 years the sum of all their ages will be 100 years. How old are the two siblings at present?

Found 2 solutions by Mini-, MathTherapy:
Answer by Mini-(17)

Let the son is X years old.
Then the daughter is 3*X years old.
The sum of all three persons ages in 10 years will be:
100 = X+10 + 3*X+10 + 42+10.
100 = 4*X + 10+10+42+10.
100 = 4*X +72.
4*X = 100-72 = 28.
X = 7.
The son is 7 years old and the daughter is 28 years old.

Answer by MathTherapy(10552)


A father who is 42 years old has a son and a daughter. The daughter is three times as old as the son. In 10 years the sum of all their ages will be 100 years. How old are the two siblings at present?